# Not flexible enough? Impacts of electric carsharing on a power sector with variable renewables11footnotemark: 1

### v1

@@ 3.3 Shared BEVs can still charge and discharge in line with variable renewables, but to a lesser extent
We observe that, with perfect foresight, BEV fleets with shared cars could still be operated largely in line with variable renewable generation patterns (Figures 6 and 16). In particular, optimized charging takes place when RES availability is high, e.g., during negative residual load events in summer, and during periods with low positive residual load in winter (Figure 14). Conversely, BEVs are discharged (in the case of bidirectional charging) at times when the residual load is high, meaning that the demand exceeds variable renewable electricity generation.
A fleet with shared BEVs has a lower potential to balance renewable variations, i.e., to shift energy between negative and positive residual load periods, than a fleet of only private BEVs: light and dark blue areas on Figure 6 are always larger than or equal to the sum of the purple and pink areas. The difference in magnitude is smaller in the low uptake case (Figure 15). The charging and discharging patterns of the shared fleet (light purple and light pink) are more variable compared to those of the overall fleet. The reason is that shared vehicles are driven much more frequently and hence can only charge or discharge for shorter periods before they are used again. These results also apply to the “Shared only” setting (Figure 17).
@@ 3.4 Green hydrogen production as another source of flexibility
In additional model runs, we include green hydrogen production as an example of another sector coupling technology which increases the demand-side flexibility of the system. More precisely, we consider a yearly industrial demand for green hydrogen of 30 TWh, which is assumed to be produced domestically via electrolysis. This largely reflects current German hydrogen policy targets. Note that this demand for green hydrogen comes on top of the hydrogen required for reconversion to electricity, i.e., for long-duration electricity storage. We assume that the industrial hydrogen demand is equally distributed over every hour of the year. In general, green hydrogen can be flexibly produced at low costs in periods when plenty of renewable electricity is available and prices are low. Yet, shifting hydrogen production to such periods requires according electrolysis and hydrogen storage capacities. The model optimizes these capacities as well as their hourly use as endogenous model variables, rendering hydrogen production a flexible sector coupling option.
@@
Including additional green hydrogen production in the model leads to lower overall cost increases of carsharing both in absolute and relative terms (Table 9). Additional costs per substituted car decrease by 14-22% in the “Shared + other BEVs” framework and by 18-32% in the “Shared only” framework, compared to the setting with no green hydrogen industrial demand. The reason for this is that green hydrogen production competes with smartly charged BEVs for cheap electricity in periods where renewable availability is high. Accordingly, smart and bidirectional charging are less beneficial in the reference than in the setting without hydrogen. In turn, reducing the BEV fleet’s flexibility potential by shifting to carsharing also hurts less.
This is also visible in the effects of carsharing on optimal generation and storage capacities. Irrespective of the uptake scenario, carsharing triggers lower overall investments into renewable and firm generation capacity as well as electricity storage as compared to the case without green hydrogen (Figures 26 to 29). In particular, optimal renewable capacities decrease. Besides, the reduction in wind generation capacity is even more pronounced in the case with a green hydrogen industrial demand. This is because the optimal renewable share exceeds 80% in the reference with industrial green hydrogen demand.
4 Discussion of model choices and limitations
@@
Our model results rely on a set of assumptions. The parameterization is on purpose relatively restrictive when it comes to the flexibility of the power sector: it does not include interconnection with other neighbouring countries, nor other flexible demand-side options that could engage in temporal load shifting. Including additional flexibility options would likely decrease the impacts of carsharing, which would translate into smaller cost increases as well as smaller deviations from the reference in terms of capacity mix and dispatch. In this respect, our main results can be interpreted as an upper bound of the power sector impacts.
@@
When it comes to the specific mobility time series, our framework is also rather conservative in terms of flexibility. In the “Shared only” framework, we consider only one cluster that switches to carsharing, namely the one with the smallest amount of trips per day and with the shortest overall travelled distance. These vehicles are idle most of the time, and thus offer the greatest flexibility potential if smart charging or vehicle-to-grid technologies are available. Assuming that such profiles switch entirely to carsharing removes the greatest source of flexibility that one could get from the sector coupling with passenger cars. The results thus provide a very conservative quantification of how much the flexibility potential of BEVs would be reduced. The same applies to the “Shared + other BEVs” framework, although to a smaller extent, since we also consider most intense car usage by including clusters 3 to 5. Cost increases still remain moderate.
@@
Another key assumption of our setting is that car mobility needs remain constant, irrespective of the way they are fulfilled. This means that the switch to carsharing does not induce a modification of when and how people decide to move with a car. This is disputable as it could be the case that switching to carsharing comes along with less car trips or shorter car trips, potentially combined with other transport modes such as active modes or public transport. In the case where carsharing would induce a decrease in overall car use, this could entail a greater flexibility of the shared cars than the one we model in this paper, for a constant size of the shared BEV fleet. This corroborates our upper bound interpretation of the modeled flexibility reduction effects.
@@
Next, car users and fleet operators are assumed to have perfect foresight when it comes to charging and discharging their cars. This means that people can perfectly predict their car trips over time. This absence of uncertainty when it comes to the timing and duration of future trips enables to optimize the grid interactions of the overall fleet to the fullest extent possible, providing a benchmark that may not be achievable with more realistic behaviours. In that sense, our modelling might overestimate the flexibility of both private and shared BEVs. Note that we focus on deviations to respective references, such that assumptions present in both settings play less of a role.
@@
Some modelling features rely on stylized assumptions for tractability and numerical feasibility, especially concerning BEVs time series. For example, including more car mobility profiles and additional vehicle models would be desirable to generate an even more representative fleet, but this heavily increases the numerical burden. Likewise, we assume that shared cars always find a charging station after a trip and that such charging stations have a power rating of 75 kW. Even if this is optimistic compared to today’s situation, it appears plausible that future carsharing fleets could be managed by their operators such that users can always find a charging station, potentially a specifically designed one for carsharing users.
On the power modelling side, a potential limitation is the optimization at a national scale and not at the substation level, which could potentially overestimate the flexibility potential of BEVs with smart charging (strobel2022joint (38)).
5 Conclusion
The results indicate that a switch to electric carsharing decreases the flexibility potential of BEVs and thus increases electricity sector costs. This cost increase however remains moderate in most settings, and remains below 100 euros per substituted car and per year in our main scenario. Cost increases are highest when other sources of flexibility are not available, such as private BEVs or green hydrogen production, and when all electric vehicles are assumed to be charged bidirectionally, i.e., make full use of their flexibility potential.
Costs effects of carsharing are driven by changes in the optimal power plant portfolio and its hourly use, which also sheds light on the benefits of coupling a largely renewable power sector with BEVs. In general, switching to carsharing increases the optimal share of solar PV and in turn decreases the share of wind power, as the aggregate vehicle battery storage capacity decreases. Effects on stationary electricity storage needs depend on various factors. In our main scenario, we find that the increase in stationary storage caused by carsharing is much lower than the substituted battery capacity of privately owned cars. Concerning the use of mineral resources for batteries, both for vehicles and stationary storage, carsharing is thus likely to have an overall relieving effect.
Overall, our research demonstrates that future BEV fleets that include relevant amounts of shared cars can still be operated largely in line with variable renewable energy sources and thus contribute to their system integration. Note that this relies on the assumption that vehicle operators, or some aggregators, have perfect foresight on mobility patterns and charging availability, both for privately-owned and shared BEVs. Hence, our results imply an upper bound of the beneficial system effects of flexible BEVs. Yet, we argue that system-oriented charging of future shared car fleets appears to be plausible, given that carsharing operators are likely to have an interest in low charging costs.
@@ 5 Conclusion
While most effects determined in our analysis are relatively small in size, both in terms of cost increases and changes in the capacity mix, they should be put in perspective with potential gains from switching to electric carsharing and reducing the BEV fleet size. BEVs are also associated with negative externalities which can be thought as social costs (gossling2019social (15, 7)). Methods to derive estimates of such costs have been developed (schroder2023ending (37)). In a case study for Munich, the overall external cost of a BEV is estimated to be around 14 euro-cents per kilometer driven. Assuming BEVs are driven ten kilometers a day, and five days per week, the yearly external cost of a BEV is around 364 euros. Removing one BEV from the system could thus represent a gain of 364 euros per year for society. In our main scenario, these reduced negative externalities are around four times larger than power sector cost increases under the assumption of bidirectional charging, and even around a factor eight larger assuming smart charging. These estimates need to be interpreted cautiously since they depend on many underlying assumptions and are specific to each urban setting. Nevertheless, they highlight that the increase in the power sector costs is only a part of the overall impact of switching from private to shared cars. In addition, carsharing may also reduce the carbon footprint of BEVs because of higher overall use of cars over their lifetime (morfeldt2022 (30)). We conclude that the potential societal gains from the reduction of the car fleet size are likely to outweigh the potentially adverse electricity sector effects. In this light, incentives and support to implement effective and profitable sharing services appear meaningful along with further research on how to most optimally design and operate such fleets and the charging infrastructure supporting them.
@@
Adeline Guéret: Conceptualization, Methodology, Software, Formal analysis, Investigation, Data Curation, Writing - Original Draft, Writing - Review & Editing, Visualization. Wolf-Peter Schill: Conceptualization, Investigation, Writing - Review & Editing, Funding acquisition. Carlos Gaete-Morales: Software.
@@ Acknowledgments
[…]
@@ Appendix A Clusters
The driving electricity consumption time series computes for 5-minutes time step how much electricity is used to power the trip depending on vehicle and climatic features such as the motor or the ambient temperature as well as the speed of the trip.
The grid availability time series is also computed at the 5-minutes time step and indicates whether the vehicle is plugged to a charging station and if yes, at which power rating.
@@ Appendix C DIETER
Hard coal 0.0 9.8 0.04 35 0.955 3845 30 0.430 0.337 8.27
Lignite 0.0 9.3 0.04 35 0.953 3845 30 0.380 0.399 5.50
Natural gas (OCGT) 0.0 +Inf 0.04 25 0.948 435 15 0.400 0.201 30.00
Natural gas (CCGT) 0.0 +Inf 0.04 25 0.960 830 20 0.542 0.201 30.00
Bio energy 6.0 6.0 0.04 30 1.0 3300 100 0.487 0.000 32.50
Run-of-river 5.6 5.6 0.04 50 1.0 3312 30 0.900 0.000 0.00
Photovoltaic 0.0 +Inf 0.04 25 1.0 519 25 1.000 0.000 0.00
Onshore wind 0.0 +Inf 0.04 25 1.0 1035 35 1.000 0.000 0.00
Offshore wind 0.0 +Inf 0.04 25 1.0 1934 100 1.000 0.000 0.00
@@ D.1 Power sector costs
Reference (in billion euros) 56.3 57.6 55.6 56.2 54.4 54.8
Difference (in million euros) 21.6 8.2 196.5 429.4 977.5 1214.0
[…]
Reference (in billion euros) 58.3 59.6 57.7 58.4 56.7 57.3
Difference (in million euros) 19.7 14.3 160.4 295.4 734.3 847.6
[…]
@@
Reference (in billion euros) 61.5 61.5 58.4 58.4 57.0 57.0
Difference (in million euros) 6.6 -15.2 163.4 338.3 355.0 613.3
[…]
Reference (in billion euros) 63.5 63.5 60.9 60.9 59.7 59.7
Difference (in million euros) 5.3 -12.1 132.7 263.9 302.5 518.9
[…]
@@
D.2 Power generation and capacity mix

### v2

@@ 3.3 Shared BEVs can still charge and discharge in line with variable renewables, but to a lesser extent
We observe that, with perfect foresight, BEV fleets with shared cars could still be operated largely in line with variable renewable generation patterns (Figures 6 and 18). In particular, optimised charging takes place when RES availability is high, e.g. during negative residual load events in summer, and during periods with low positive residual load in winter (Figure 16). Conversely, BEVs are discharged at times when the residual load is high, meaning that the demand exceeds variable renewable electricity generation.
A fleet with shared BEVs has a lower potential to balance renewable variations, i.e. to shift energy between negative and positive residual load periods, than a fleet of only private BEVs: light or dark blue areas on Figure 6 are always larger than or equal to the sum of the purple and pink areas. The difference in magnitude is slightly smaller in the low uptake case (Figure 17). The charging and discharging patterns of the shared fleet (light purple and light pink) are more variable compared to those of the overall fleet. The reason is that shared vehicles are driven much more frequently and hence can only charge or discharge for shorter periods before they are used again. These results also apply to the “Shared only” setting (Figure 19).
@@ 3.4 Green hydrogen production as another source of flexibility
In additional model runs, we include green hydrogen production as an example of another sector coupling technology which increases the demand-side flexibility of the system. More precisely, we consider a yearly industrial demand for green hydrogen of 30 TWh, which is assumed to be produced domestically via electrolysis. This largely reflects current German hydrogen policy targets for 2030. We assume that this demand for green hydrogen comes on top of the hydrogen required for reconversion to electricity, i.e. for long-duration electricity storage, and that the industrial hydrogen demand is equally distributed over every hour of the year. In general, green hydrogen can be flexibly produced at low costs in periods when plenty of renewable electricity is available and prices are low. Yet, shifting hydrogen production to such periods requires according electrolysis and hydrogen storage capacities. The model optimises these capacities as well as their hourly use as endogenous model variables, rendering hydrogen production a flexible sector coupling option.
@@
Including additional green hydrogen production in the model does not lead to major changes of the power sector cost effects of carsharing (Table 9). In the “Shared + other BEVs” framework, the additional costs per substituted car slightly decrease by 4-6%, compared to the setting with no green hydrogen industrial demand. The reason for this is that green hydrogen production serves as a flexible load, which makes smart and bidirectional BEV charging less beneficial than in the setting without hydrogen. Reducing the BEV fleet’s flexibility potential by shifting to carsharing thus hurts less. In the “Shared only” framework, relative effects are even smaller. One of the reasons why the power sector cost effects of carsharing are robust against adding flexible green hydrogen production is that the renewable share slightly differs (Table 10). Especially in the cases with bidirectional charging, carsharing induces a slightly higher use of non-renewable electricity generation when green hydrogen production is added compared to the baseline specification, which helps to reduce power sector costs effects. This goes along with respective changes in renewable generation capacity (Figure 28 and Figure 29). In alternative model parameterisations where the minimum renewable share is always binding, we conversely find that green hydrogen production increases the power sector costs effects of carsharing, as electrolysis would then compete with smartly charged BEVs for cheap electricity in periods where renewable availability is high.
4 Discussion
@@
Our model results rely on a set of assumptions. The parameterisation is on purpose rather restrictive when it comes to the flexibility of the power sector: it does not include interconnection with other neighbouring countries, nor other flexible demand-side options that could engage in temporal load shifting. Including additional flexibility options would likely decrease the impacts of carsharing, which would translate into smaller cost increases as well as smaller deviations from the reference in terms of capacity mix and dispatch. In this respect, our main results can be interpreted as an upper bound of the power sector impacts.
@@
When it comes to the specific mobility time series, our framework is also rather conservative in terms of flexibility. In the “Shared only” framework, we consider only one cluster that switches to carsharing, namely the one with the smallest amount of trips per day and with the shortest overall travelled distance. These vehicles are idle most of the time, and thus offer the greatest flexibility potential if smart charging or vehicle-to-grid technologies are available. Assuming that such profiles switch entirely to carsharing removes the greatest source of flexibility that one could get from the sector coupling with passenger cars. The results thus provide a very conservative quantification of how much the flexibility potential of BEVs would be reduced. The same applies to the “Shared + other BEVs” framework, although to a smaller extent, since we also consider more intense car usage by including clusters 3 to 5. Cost increases still remain moderate.
@@
Another key assumption of our setting is that car mobility needs remain constant, irrespective of the way they are fulfilled. This means that the switch to carsharing does not induce a modification of when and how people decide to move with a car. This is disputable as it could be the case that switching to carsharing comes along with less car trips or shorter car trips, potentially combined with other transport modes such as active modes or public transport. If carsharing induced a decrease in overall car use, this could entail a greater flexibility of the shared cars than the one we model in this paper, hence a smaller cost increase. This is confirmed by a sensitivity analysis where we assume shared cars are used less, such that their electricity consumption decreases by 10% (see Table 11 and Figures 44 and 45), which corroborates our upper bound interpretation of the modelled flexibility reduction effects.
Previous research has shown that carsharing has the potential to decrease private car ownership, and one shared car might replace up to 20 private cars in the most optimistic cases ([jochem_does_2020]). Nevertheless, whether carsharing ultimately reduces car ownership and contributes to a decrease of the fleet size relies on an efficient policy design without which carsharing might even lead to an increase of the car fleet size. For example, carsharing policies should ensure that carsharing is used as a complement rather than a substitute to public transport and active mobility ([silvestri2021contribution]). In this work, we assume a substitution rate of about five, which reflects a decrease in private car ownership, but remains conservative in comparison to some potential substitution rates found in the literature. Assuming a higher substitution rate might lead to higher cost increases, but it could also imply greater mobility changes, such as smaller and/or less numerous car trips. More research needs to be carried out in order to better understand how these effects interact.
@@
Next, car users and fleet operators are assumed to have perfect foresight when it comes to charging and discharging their cars. This means that people can perfectly predict their car trips over time. This absence of uncertainty when it comes to the timing and duration of future trips enables to optimize the grid interactions of the overall fleet to the fullest extent possible, providing a benchmark that may not be achievable with more realistic behaviours. In that sense, our modelling might overestimate the flexibility of both private and shared BEVs. Note that we focus on deviations to respective references, such that assumptions present in both settings play less of a role. Besides, we consider that the whole BEV fleet uses the same charging strategy. This might lead to an overestimation of the flexibility potential when smart or bidirectional charging is considered. Nevertheless, considering only one charging strategy across the fleet can be interpreted as a limit case. Results for mixed fleets with several charging strategies should lie in between these limit cases. Integrating mixed fleets in modelling exercises would allow to better represent the different abilities of carsharing operators and private car owners to choose flexible (dis-)charging strategies and could be done in further research.
@@
Some modelling features rely on stylised assumptions for tractability and numerical feasibility, especially concerning BEVs time series. For example, including more car mobility profiles and additional vehicle models would be desirable to generate an even more representative fleet, but this heavily increases the numerical burden. On the power modelling side, a potential limitation is the optimisation at a national scale and not at the substation level, which could potentially overestimate the flexibility potential of BEVs with smart charging ([strobel2022joint]). Further research on how to increase the representativity of mobility patterns while not under- or overestimating the flexibility potential of BEVs in power sector models is still needed. Besides, we assume a bigger battery capacity for shared BEVs compared to private ones. Our assumption is also slightly bigger than what is available on the carsharing market today, but it appears plausible that this will not be the case in the future anymore. Assuming a smaller battery capacity for shared cars would further constrain the flexibility potential of shared BEVs, which might lead to higher cost increases. Nevertheless, too small a battery size would not allow vehicles to accommodate as many trips as required by the constant mobility need assumption. In this regard, it seems that the overestimation of the flexibility potential of shared cars remains small. Likewise, we assume that shared cars always find a charging station after a trip and that such charging stations have a power rating of 75 kW. Even if this is optimistic compared to today’s situation, it appears plausible that future carsharing fleets could be managed by their operators such that users can always find a charging station, potentially a specifically designed one for carsharing users.
@@
While most effects determined in our analysis are relatively small in size, both in terms of cost increases and changes in the capacity mix, they should be put in perspective with potential gains from switching to electric carsharing and reducing the BEV fleet size. BEVs are also associated with negative externalities which can be thought as social costs ([gossling2019social, de2018total]). Methods to derive estimates of such costs have been developed ([schroder2023ending]). In a case study for Munich, the overall external cost of a BEV is estimated to be around 14 euro-cents per kilometer driven. Assuming BEVs are driven ten kilometers a day, and five days per week, the yearly external cost of a BEV is around 364 euros. Removing one BEV from the system could thus represent a gain of 364 euros per year for society. In our main scenario, these reduced negative externalities are around three times larger than power sector cost increases under the assumption of bidirectional charging, and even around a factor seven larger assuming smart charging. These estimates need to be interpreted cautiously since they depend on many underlying assumptions and are specific to each urban setting. Nevertheless, they highlight that the increase in the power sector costs is only a part of the overall impact of switching from private to shared electric cars. In addition, carsharing may also reduce the carbon footprint of BEVs because of higher overall use of cars over their lifetime ([morfeldt2022]). An evaluation of the extent to which the changes in power sector carbon emissions might be counteracted by the reduction in emissions related to car and battery manufacturing could constitute future research work. From these insights, we conclude that the potential societal gains from the reduction of the car fleet size are likely to outweigh the potentially adverse electricity sector effects. Nevertheless, more research needs to be done in this direction, in order to better understand how different effects play one with another. Further interdisciplinary research on how to optimally design and operate carsharing fleets and the charging infrastructure supporting them is also required.
Overall, our research demonstrates that future BEV fleets that include relevant amounts of shared cars can still be operated largely in line with variable renewable energy sources and thus contribute to their system integration. Note that this relies on the assumption that vehicle operators, or some aggregators, have perfect foresight on mobility patterns and charging availability, both for privately owned and shared BEVs. Hence, our results imply an upper bound of the beneficial system effects of flexible BEVs. Yet, we argue that system-oriented charging of future shared car fleets appears to be plausible, given that carsharing operators are likely to have an interest in low charging costs.
We conclude that the moderate power system cost increases related to electric carsharing are only one aspect of various, and potentially beneficial, changes brought by such a structural shift in private transport. Further research has to be done to comprehensively evaluate how all these aspects interact one with another. At the same time, policymakers should strive to enhance the ability of carsharing operators to operate their fleet as flexibility as possible and design policies ensuring that carsharing enables a decrease in private car ownership.
@@
Adeline Guéret: Conceptualisation, Methodology, Software, Formal analysis, Investigation, Data Curation, Writing - Original Draft, Writing - Review & Editing, Visualisation. Wolf-Peter Schill: Conceptualisation, Investigation, Writing - Review & Editing, Funding acquisition. Carlos Gaete-Morales: Software, Visualisation.
Data and code availability
@@ Acknowledgments
We thank the members of the research group Transformation of the Energy Economy at DIW Berlin and participants to the ITEA Conference 2023, the European Transport Conference 2023, the British Institute of Energy Economics Research Conference 2023, the YEEES 32 Seminar 2023, the International Energy Workshop 2024, the EAERE Conference 2024, the 5th CIRED Summer School and the Chalmers Transport Seminar, as well as participants to the DIW EVU/KLI Brown Bag Seminar and the DIW Graduate Center Workshop 2023 for their constructive feedback and helpful comments. We acknowledge research grants by the German Federal Ministry of Education and Research via the “Ariadne” projects (Fkz 03SFK5NO & 03SFK5NO-2).
Declaration of interests
The authors declare no competing interests.
@@ Appendix A Clusters
The driving electricity consumption time series computes for 5-minutes time steps how much electricity is used to power the trip depending on vehicle and climatic features such as the motor or the ambient temperature as well as the speed of the trip.
The grid availability time series indicates whether the vehicle is plugged to a charging station and if yes, at which power rating. The grid availability time series is also computed at the 5-minute time step and depends on the assumed average charging station availability (Table 5) and the mobility pattern of a particular vehicle. For shared BEVs, we assume a higher charging station availability than for private BEVs, as we assume that carsharing operators are able to design optimally both users’ financial incentives and their charging network so that customers can always plug the shared BEV they have used after a ride.
@@
Figure 9: Density of number of daily trips by location and day type in the reference and in the scenario [“Shared + other BEVs”, high uptake]
Figure 10: Distribution of BEVs’s parking place for each time step by location and day type in the reference and in the scenario [“Shared + other BEVs”, high uptake]
@@ Appendix C DIETER
Hard coal 0.0 9.8 0.04 40 0.955 3845 61.520 0.430 0.337 8.27
Lignite 0.0 9.3 0.04 40 0.953 3845 61.520 0.380 0.399 5.50
Natural gas (OCGT) 0.0 +Inf 0.04 25 0.948 435 7.830 0.400 0.201 30.00
Natural gas (CCGT) 0.0 +Inf 0.04 25 0.960 830 27.390 0.542 0.201 30.00
Bio energy 6.0 6.0 0.04 25 1.0 3300 118.80 0.487 0.000 32.50
Run-of-river 4.73 4.73 0.04 80 1.0 3312 66.240 0.900 0.000 0.00
Photovoltaic 0.0 +Inf 0.04 30 1.0 519 12.975 1.000 0.000 0.00
Onshore wind 0.0 +Inf 0.04 27 1.0 1035 13.455 1.000 0.000 0.00
Offshore wind 0.0 +Inf 0.04 27 1.0 1934 36.746 1.000 0.000 0.00
@@ D.1 Power sector costs
Reference (in billion euros) 44.7 45.7 44.1 44.5 43.2 43.5
Difference (in million euros) 27.0 81.6 105.0 282.5 672.8 939.3
No H2 demand Difference (in euros per substituted car) 6.7 13.1 27.1 45.2 173.6 150.3
Reference (in billion euros) 46.0 47.0 45.4 45.9 44.5 44.8
Difference (in million euros) 25.4 80.1 108.5 284.5 674.4 928.4
Green H2 demand Difference (in euros per substituted car) 6.5 12.8 28.0 45.5 174.1 148.6
@@
Reference (in billion euros) 48.9 48.9 46.0 46.0 44.9 44.9
Difference (in million euros) -86.0 -45.3 134.4 306.6 382.5 676.4
No H2 demand Difference (in euros per substituted car) -22.3 -7.3 34.7 49.2 98.9 108.4
Reference (in billion euros) 50.2 50.2 47.4 47.4 46.3 46.3
Difference (in million euros) -85.9 -46.6 126.2 293.2 357.8 640.7
Green H2 demand Difference (in euros per substituted car) -22.2 -7.5 32.6 47.0 92.5 102.7
@@
D.2 Renewable energy shares
Shared only
Uncontrolled Smart charging Bidirectional
Low High Low High Low High
Reference (in %) 85.5 85.5 85.5 85.6 86.6 87.2
No H2 demand Difference (in %) 0.1 0.1 -0.1 0.0 -0.9 -1.3
Reference (in %) 85.2 85.0 85.2 85.6 86.9 87.8
Green H2 demand Difference (in %) 0.1 0.3 0.0 -0.1 -1.5 -1.8
Shared + other BEVs
Uncontrolled Smart charging Bidirectional
Low High Low High Low High
Reference (in %) 85.2 85.2 86.5 86.5 88.0 88.0
No H2 demand Difference (in %) 0.1 0.3 0.0 -0.3 -0.2 -0.6
Reference (in %) 84.8 84.8 86.9 86.9 89.5 89.5
Green H2 demand Difference (in %) 0.2 0.2 0.0 0.0 -0.8 -1.4
Table 10: Share of renewable energy in electricity consumption for reference and scenarios across frameworks, uptake regimes and charging strategies
D.3 Power generation and capacity mix